I need assistance in writing the slope-intercept equation for the line with the given slope that contains the given point.
m = 3; (6, 2)
Thank you so much!

Answer by sudhanshu_kmr(1152)


equation is (y-2) = 3(x-6)
=> y-2 = 3x -18
=> y = 3x -18 +2
=> y = 3x -16

now, equation is in slope intercept form i.e y = mx + b
where m= 3 and b = -16